OpenStudy (anonymous):

Why would knowing the Pythagorean identity sin^2x+cos^2x=1 and the sum identities allow you to prove all the other identities you have seen in this lesson?

OpenStudy (princeharryyy):

'cos sinx and cosx are the only two functions that exists in reality ..... so everything else is a combination of sinx and cosx @MrClaymoar

jhonyy9 (jhonyy9):

hope you know a^2 +b^2 =c^2 what is this ?

OpenStudy (princeharryyy):

i thik that's pythagoras theorem for any triangle that has one angle as 90 deg... @jhonyy9

jhonyy9 (jhonyy9):

and how do you think it this not is same sin^2 x +cos^ 2 x = 1^2 ??? not are the same ?